Martin Company purchases a machine at the beginning of the year at a cost of $80,000. The machine is depreciated using the double-declining-balance method. The machine’s useful life is estimated to be 4 years with a $6,600 salvage value. The machine’s book value at the end of year 3 is:
Answer #1

Double declining rate = 200/useful life = 200/4 = 50%

Depreciation for year 1 = 80,000 * 50% = 40,000

Depreciation for year 2 = (80,000-40,000)*50% = 20,000

Depreciation for year 3 = (40,000-20,000) * 50% = 10,000

Book value at the end of year 3 = 80,000 - (40,000+20,000+10,000)

= 10,000